If the difference of a number and four is doubled, the result is 1/4 less than the number. Find the number.
Found 2 solutions by Adam, edjones:
Answer by Adam(64)

We denote the number x
difference of a number and four (x-4)
is doubled 2*(x-4)
he result is 1/4 less than the number 2*(x-4)=x-1/4
2x-8=x-1/4
2x-x=8-1/4
x = 8-1/4
x = 7 3/4

Answer by edjones(8007)

2(x-4)=3x/4
2x-8=3x/4
multiply 4 times both sides 8x-32=3x
subtract 3x and add 32 to both sides: 5x=32
divide 5 into each side: x=32/5
Check:
(32/5)-(20/5)=12/5
(12/5)*2=24/5
(32/5)*(1/4)=8/5
32/5-8/5=24/5
24/5=24/5
Ed